What is the mass of water required to fill a circular hot tub 2.4 meters in diameter and 1.2 meters deep answer in units of kilograms?
1
Expert's answer
2013-09-02T10:24:18-0400
m = ρ* V = ρ * S * H = ρ * Pi * (d^2) / 4 * H
Where:
ρ - density of water [ =
1000 kg / (m^3) ]
V - volume of circular hot tub
S - area of bases of circular hot tub [ = Pi * (d^2 / 4) ]
Pi ≈3.14
H = 1.2 m
d = 2.4 m
So,
m ≈ 1000 * 3,14 * (2.4^2) * 1.2 / 4 ≈ 5425,92
(kg)

ANSWER: m ≈ 5425,92 (kg)